# Explanation User Interfaces: A Systematic Literature Review

### v1

@@ 5.1. Relationships Between AI Systems and Explainability Techniques
The distribution of XAI methodologies, as shown in Table 7, reveals that feature importance methods are the most frequently used, followed by counterfactual explanations and Shapley values. These techniques primarily aim to provide localised insights into individual model predictions. The dominance of neural networks and the frequent use of these explanation techniques underscore the need for some form of interpretability in such complex systems. Feature importance, counterfactual explanations, and Shapley values are valuable tools to bridge the gap between black-box models and actionable, user-friendly explanations. As emphasised by Miller, effective explanations should focus on relevance and comprehensibility rather than exhaustively detailing model structure (Miller, 2019). It should still be noted that, although the literature employs such techniques to provide some form of explanation for the black-box model, this has a polarising effect as some argue that this is, in fact, not the best course of action, instead preferring white-box models when possible (Rudin, 2019).
@@
Partial Dependency Plot 4
@@
Some less common techniques, such as salient masks (Hwang et al., 2022a; Kadir et al., 2023), neuron activation (Wang et al., 2023), decision trees (Khanna et al., 2022), and partial dependency plots, suggest niche or specialised applications in XUIs. These methods often target unique requirements or particular user groups, emphasizing the diverse needs within the XUIs design space.
@@
With regard to the relationship between XAI techniques and AI algorithms, several recurring patterns can be observed. Feature importance is prominently linked to neural networks (Lee et al., 2020; Hernandez-Bocanegra and Ziegler, 2021; Röhrl et al., 2023; Lee et al., 2023; Heimerl et al., 2022; Mishra et al., 2022; Shin et al., 2022), ensembles (Malandri et al., 2023; Bhattacharya et al., 2024; He et al., 2024), and probabilistic methods (Riveiro and Thill, 2021), reflecting its adaptability across diverse AI models. Counterfactual explanations are broadly applied, particularly in ensemble (Malandri et al., 2023; Ma et al., 2022; Bove et al., 2023; Cau et al., 2023) and neural network settings (Sovrano and Vitali, 2021; Esfahani et al., 2024; Zhang et al., 2024; Hao et al., 2023). Shapley values are closely associated with neural networks (Lu et al., 2024; Kim et al., 2023c; Meza Martínez et al., 2023; Wang et al., 2022; He et al., 2024) and probabilistic models (Malandri et al., 2023; Purificato et al., 2023; Cheng et al., 2022), leveraging their utility in quantifying feature contributions in complex systems. Techniques like neuron activation are nearly exclusive to neural networks (Hwang et al., 2022a; Wang et al., 2023; Nourani et al., 2020; Ahn et al., 2024; Wu et al., 2021) due to their specific relevance to such architectures. Decision rules appear with rule-based (Piorkowski et al., 2023; Sevastjanova et al., 2021), ensemble (Bhattacharya et al., 2024; Wang et al., 2019) and Neural Network (Ming et al., 2019; Meza Martínez et al., 2023; Krause et al., 2017), while decision trees are used along with reinforcement learning techniques (Khanna et al., 2022; Kridalukmana et al., 2022).
@@
Overall, the choice of XAI techniques mostly depends on the type of AI algorithm being explained. Certain XAI techniques are designed for specific model types, such as decision trees in reinforcement learning, neural activation analysis for neural networks, and rule extraction for rule-based models. In contrast, more versatile methods, including feature importance, Shapley values, and counterfactual examples, are widely adopted due to their adaptability across various contexts and model architectures.
5.2. Explanation Modalities
The presentation of explanations to users is crucial in determining the effectiveness and usability of XUIs. Table 8 categorises the explanation modalities observed in the surveyed studies. Visual explanations dominate, likely due to their intuitive and accessible nature. Formats such as graphs, heatmaps, and diagrams translate complex AI reasoning into interpretable insights for users. Text-based explanations are also widely adopted, reflecting the value of concise descriptions for clarity.
Figure 3. Frequency of visualisation techniques used in explanation user interfaces. The bar chart presents the types of visualisations employed to convey explanations, with heatmaps, bar charts, and trend lines being the most frequently used. A wide range of other formats—such as scatter plots, histograms, Sankey diagrams, and tree structures—are also represented, reflecting the diversity of visual explanation strategies.
[…]
Frequency of visualisation techniques used in explanation user interfaces. The bar chart presents the types of visualisations employed to convey explanations, with heatmaps, bar charts, and trend lines being the most frequently used. A wide range of other formats—such as scatter plots, histograms, Sankey diagrams, and tree structures—are also represented, reflecting the diversity of visual explanation strategies.
Natural language explanations, audio, and tangible formats are comparatively underutilised, potentially reflecting their more specialised use cases or higher development costs. Video explanations are rare, possibly due to the time-intensive nature of their production and the immediate interactivity required in many XUIs.
@@ 5.2. Explanation Modalities
Figure 3 details the visualization types employed in XUIs. Heatmaps stand out as the most frequently used visualization, likely due to their ability to quickly convey patterns or feature attributions. Bar charts and trend lines are common, reflecting their utility in comparative and temporal data analysis. Hybrid approaches that integrate textual and visual elements seem to be preferred in the literature, suggesting that combining modalities may provide the most effective explanations by complementing graphical clarity with textual context.
Table 8. Count of explanation modalities used in XUIs.
Explanation Modality Count
@@
5.3. Interactivity
Interactive XUIs allow users to manipulate explanation parameters, explore different perspectives, or request additional details, potentially enhancing comprehension and trust (Weitz et al., 2021; Gromowski et al., 2020). These interfaces often employ dynamic visualization techniques, adjustable feature importance rankings, or interactive counterfactual explanations, enabling users to tailor explanations to their needs.
Conversely, non-interactive XUIs present static explanations without user input (Riveiro and Thill, 2021; Upasane et al., 2021). While these may still provide valuable insights, they often lack adaptability, potentially limiting their effectiveness in complex decision-making scenarios. The prevalence of interactive systems in the dataset suggests a growing emphasis on user engagement in XAI research, highlighting the importance of designing explanations that support exploratory and user-driven interactions.
6. RQ3. What solutions are used to evaluate XUIs?
This research question explores the relationship between the user, the type of user study conducted, and the metrics employed, bridging the Human-Computer Interaction perspective with the algorithmic approach to offer a comprehensive understanding of the XUIs environment. In the development of XUIs, evaluating how users interact with and perceive these systems is critical to ensuring their effectiveness, usability, and trustworthiness (Abdul et al., 2018). The way users are evaluated and the metrics observed play a central role in understanding whether the explanations provided by an AI system align with user needs and expectations. By examining study methodologies and key performance metrics, researchers can identify strengths and gaps in XUI design, ensuring that these systems are technically robust and user-centered. This analysis is essential for advancing the field of XAI and fostering greater adoption of AI systems in real-world applications where transparency and trust are paramount.
@@ 5.3. Interactivity
Table 10. Top 14 Evaluation Metrics for User study
Evaluation Metrics Count
@@
Perceived Effectiveness 14
Task Performance 14
@@
6.1. Evaluation Methods and Metrics in XUI Research
@@
Table 10 categorises the various study methods used to evaluate XUIs. Controlled experiments are the most frequent ones, highlighting the importance of rigorous, systematic evaluation under controlled conditions. This method allows researchers to isolate specific factors and measure their impact on user interaction with XUIs. Additionally, interviews and usability studies are frequently utilised, highlighting the focus on user perspectives and ensuring that the interfaces remain practical and accessible.
@@
Moving onto the metrics, illustrated in Table 10, trust stands out as a central metric, indicating that fostering user confidence in AI systems is a key priority for XUI developers. Usability and workload are close behind, emphasizing the importance of ensuring the interface is easy to use and does not overwhelm the user. Perceived effectiveness and satisfaction suggest an interest in understanding users’ perceptions of how well the system supports their needs. Table 11 shows the most popular metrics. Collectively, these metrics underscore the importance of designing XUIs that not only enhance explainability but also encourage practical adoption in real-world applications. XUIs should, in fact, be designed to inspire trust, remain user-friendly, and avoid overwhelming users with excessive complexity.
@@
It is worth noting that, despite the importance of transparency in explainability research, this metric has only been explicitly assessed in eight studies, suggesting that current research in XUIs should place greater emphasis on incorporating transparency metrics into evaluations.
@@
Table 11. Papers that report user studies measuring Trust, Usability, and Workload
Metrics References
@@
Trust (Weitz et al., 2021; Kim et al., 2023d; Wysocki et al., 2023; Guo et al., 2022; Khurana et al., 2021; Lee et al., 2020; Panigutti et al., 2023; Lu et al., 2024; Karran et al., 2022; Bhattacharya et al., 2023; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Bhattacharya et al., 2024; Kim et al., 2023a; Röhrl et al., 2023; Guesmi et al., 2022; Joshi et al., 2024; Lombardi et al., 2024; Naiseh et al., 2023; Bahel et al., 2024; Guesmi et al., 2024; Lee et al., 2023; Perlmutter et al., 2024; Bertrand et al., 2023a; Park et al., 2022; Zytek et al., 2022; Purificato et al., 2023; Heimerl et al., 2022; Nimmo et al., 2024; Lei et al., 2024; Desolda et al., 2023; Stites et al., 2021)
@@
Usability (Fujiwara et al., 2020; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Wang et al., 2021; Huang et al., 2022; Cheng et al., 2021; Kim et al., 2023b; Szymanski et al., 2024; Hohman et al., 2019; Guesmi et al., 2024; Wang et al., 2022; Ming et al., 2019; Sevastjanova et al., 2021; Park et al., 2022; Zhang et al., 2024; El-Zanfaly et al., 2023; Sciascio et al., 2017; Purificato et al., 2023; Hao et al., 2023; Yuan et al., 2024b; Antar et al., 2024; Szymanski et al., 2021)
@@
Workload (Fujiwara et al., 2020; Piorkowski et al., 2023; Dominguez et al., 2020; Sanneman and Shah, 2022a; Lee et al., 2020; Wang et al., 2023; Karran et al., 2022; Khanna et al., 2022; Anderson et al., 2020; El-Zanfaly et al., 2023; Kridalukmana et al., 2022; Sciascio et al., 2017; Nakao et al., 2022; Heimerl et al., 2022; Yuan et al., 2024a)
@@
Figure 4. Methods used to evaluate explanation interfaces across different user types. The heatmap shows the distribution of evaluation methods such as controlled experiments, interviews, surveys, and usability studies, categorised by the type of user involved (e.g., AI experts, domain experts, non-experts). Controlled experiments and interviews are the most common methods, particularly among non-experts and domain experts, reflecting a strong emphasis on both empirical validation and qualitative feedback in XUIs evaluation.
[…]
@@
6.2. Tailoring XUI Evaluation Methods to User Groups
Figure 4 represents the distribution of user types across various study methodologies used in the evaluation of XUIs. The figure provides some insights into how different user groups are targeted and about the methodological approaches used to assess their interaction with and understanding of the interface. The choice of study method is influenced by the target user group: controlled experiments are favoured for domain experts and lay users due to their rigorous structure. Interviews, however, are particularly prominent in studies involving domain experts, as they can usually capture qualitative insights. Domain experts are typically involved in interviews for the application of XUI technologies in the high stakes domains such as Health (Hwang et al., 2022a; Kim et al., 2023d, c; Bhattacharya et al., 2023, 2024; Naiseh et al., 2023; Okolo et al., 2024; Larasati et al., 2023; Lindvall et al., 2021; Zhang et al., 2024; Cheng et al., 2022) and Finance and economics, (Füßl et al., 2024; Purificato et al., 2023) but also in other domain such as Education (Lu et al., 2024), Natural Language Processing (Wang et al., 2022) and weather forecasting (Kim et al., 2023a). Methods such as interactive feedback sessions, co-design, usability studies, and, notably, surveys are also employed to evaluate AI experts in various domains, such as health (Hohman et al., 2019) associated with Artificial Intelligence (Baniecki et al., 2024), education (Shin et al., 2022; Yuan et al., 2024b) and finance and economics (Hohman et al., 2019; Meza Martínez et al., 2023). While these approaches are often associated with assessing non-expert users, their application to AI experts suggests an interest in gathering qualitative and usability-focused insights from technically proficient individuals.

### v2

@@ 4.2.1. Relationships Between AI Systems and Explainability Techniques
Table 7 summarises the AI algorithms employed in the surveyed studies. Neural networks prevail over the other techniques, followed by ensemble methods and white-box models. These results highlight the prevalence of complex, opaque models like neural networks, which require XAI techniques to enhance interpretability. The distribution of XAI methodologies, as shown in Table 7, reveals that feature importance methods are the most frequently used, followed by counterfactual explanations and Shapley values. These techniques primarily aim to provide localised insights into individual model predictions. The dominance of neural networks and the frequent use of these explanation techniques underscore the need for some form of interpretability in such complex systems. Feature importance, counterfactual explanations, and Shapley values are valuable tools to bridge the gap between black-box models and actionable, user-friendly explanations. As emphasised by Miller, effective explanations should focus on relevance and comprehensibility rather than exhaustively detailing model structure (Miller, 2019). It should still be noted that, although the literature employs such techniques to provide some form of explanation for the black-box model, this has a polarising effect as some argue that this is, in fact, not the best course of action, instead preferring white-box models when possible (Rudin, 2019).
@@
Partial Dependence Plot 4
@@
Some less common techniques, such as salient masks (Hwang et al., 2022b; Kadir et al., 2023), neuron activation (Wang et al., 2023), decision trees (Khanna et al., 2022), and partial dependency plots, suggest specialised applications in XUIs. These methods often target unique requirements or particular user groups, emphasizing the diverse needs within the XUIs design space.
@@
Regarding the relationship between XAI techniques and AI algorithms, several recurring patterns can be observed. Feature importance is prominently linked to neural networks (Lee et al., 2020; Hernandez-Bocanegra and Ziegler, 2021; Röhrl et al., 2023; Lee et al., 2023; Heimerl et al., 2022; Mishra et al., 2022; Shin et al., 2022), ensembles (Malandri et al., 2023; Bhattacharya et al., 2024; He et al., 2024), and probabilistic methods (Riveiro and Thill, 2021), reflecting its adaptability across diverse AI models. Counterfactual explanations are broadly applied, particularly in ensemble (Malandri et al., 2023; Ma et al., 2022; Bove et al., 2023; Cau et al., 2023) and neural network settings (Sovrano and Vitali, 2021; Esfahani et al., 2024; Zhang et al., 2024; Hao et al., 2023). Shapley values are closely associated with neural networks (Lu et al., 2024; Kim et al., 2023c; Meza Martínez et al., 2023; Wang et al., 2022; He et al., 2024) and probabilistic models (Malandri et al., 2023; Purificato et al., 2023; Cheng et al., 2022), leveraging their utility in quantifying feature contributions in complex systems. Techniques like neuron activation are nearly exclusive to neural networks (Hwang et al., 2022b; Wang et al., 2023; Nourani et al., 2020; Ahn et al., 2024; Wu et al., 2021) due to their specific relevance to such architectures. Decision rules appear with rule-based (Piorkowski et al., 2023; Sevastjanova et al., 2021), ensemble (Bhattacharya et al., 2024; Wang et al., 2019) and Neural Network (Ming et al., 2019; Meza Martínez et al., 2023; Krause et al., 2017), while decision trees are used along with reinforcement learning techniques (Khanna et al., 2022; Kridalukmana et al., 2022).
@@
Overall, the choice of XAI techniques often depends on the type of AI algorithm being explained. While some XAI techniques are specifically designed for particular model architectures, such as neural activation analysis for neural networks (Wang et al., 2023; Hwang et al., 2022b; Nourani et al., 2020), or analyzing decision trees in reinforcement learning (Kridalukmana et al., 2022; Khanna et al., 2022). Rule extraction techniques, although associated with rule-based models (Piorkowski et al., 2023; Sevastjanova et al., 2021), can be effectively applied to a wide variety of predictive models to generate interpretable insights, regardless of the underlying algorithm (Sabbatini, 2025; Gromowski et al., 2020; Meza Martínez et al., 2023; Metta et al., 2021; Ming et al., 2019). Model-agnostic methods, on the other hand, include feature importance, Shapley values, and counterfactual examples. These models are widely adopted due to their adaptability across various contexts and model architectures (Stites et al., 2021; Riveiro and Thill, 2021; Malandri et al., 2023; Okolo et al., 2024).
4.2.2. Explanation Modalities
@@
The presentation of explanations to users is crucial in determining the effectiveness and usability of XUIs. Table 10 categorises the explanation modalities observed in the surveyed studies. Visual explanations dominate, likely due to their intuitive and accessible nature. Formats such as graphs, heatmaps, and diagrams translate complex AI reasoning into interpretable insights for users. Textual explanations are also widely adopted, reflecting the value of concise descriptions for clarity. Natural language explanations, audio, and tangible formats are comparatively underutilised, potentially reflecting their more specialised use cases or higher development costs. Video explanations are rare, possibly due to the time-intensive nature of their production and the immediate interactivity required in many XUIs. Figure 4 details the visualization types employed in XUIs. Heatmaps stand out as the most frequently used visualization, likely due to their ability to quickly convey patterns or feature attributions and their widespread use in image explanations in general. Bar charts and trend lines are common, reflecting their utility in comparative and temporal data analysis. Hybrid approaches that integrate textual and visual elements seem to be preferred in the literature, suggesting that combining modalities may provide the most effective explanations by complementing graphical clarity with textual context.
Figure 4. Frequency of visualisation techniques used in explanation user interfaces. Heatmaps, bar charts, and trend lines are the most frequently used. A wide range of other formats, such as scatter plots, histograms, Sankey diagrams, and tree structures, are also represented, reflecting the diversity of visual explanation strategies.
Frequency of visualisation techniques used in XUIs. The bar chart presents the types of visualisations employed to convey explanations, with heatmaps, bar charts, and trend lines being the most frequently used. A wide range of other formats,such as scatter plots, histograms, Sankey diagrams, and tree structures, are also represented, reflecting the diversity of visual explanation strategies.
4.2.3. Interactivity
Interactive XUIs allow users to manipulate explanation parameters, explore different perspectives, or request additional details, potentially enhancing comprehension and trust (Weitz et al., 2021; Gromowski et al., 2020). These interfaces often employ dynamic visualization techniques, adjustable feature importance rankings, or interactive counterfactual explanations, enabling users to tailor explanations to their needs. Conversely, non-interactive XUIs present static explanations without user input (Riveiro and Thill, 2021; Upasane et al., 2021). While these may still provide valuable insights, they often lack adaptability, potentially limiting their effectiveness in complex decision-making scenarios. The prevalence of interactive systems in the dataset suggests a growing emphasis on user engagement in XAI research, highlighting the importance of designing explanations that support exploratory and user-driven interactions.
4.3. RQ3. What solutions are used to evaluate XUIs?
This research question is characterized from a human-centered perspective, since it explores the relationship between the user, the type of user study conducted, and the evaluation constructs (e.g., usability, trust) employed, bridging the Human-Computer Interaction perspective with the algorithmic approach to offer a comprehensive understanding of the XUIs environment. In the development of XUIs, evaluating how users interact with and perceive these systems is critical to ensuring their effectiveness, usability, and trustworthiness (Abdul et al., 2018). The way users are evaluated and the constructs observed play a central role in understanding whether the explanations provided by an AI system align with user needs and expectations. By examining study methodologies and key performance metrics, researchers can identify strengths and gaps in XUI design, ensuring that these systems are technically robust and user-centered. This analysis is essential for advancing the field of XAI and fostering greater adoption of AI systems in real-world applications where transparency and trust are crucial.
Summary of findings
The most common evaluation methods used with XUIs are controlled experiments (typically used with domain experts and non-experts), followed by interviews (often used with domain experts, capturing qualitative data) and usablity studies. Domain experts are typically involved in evaluating XUIs in high-stakes domain (e.g., health and finance), while AI experts are also involved in usability studies and interactive feedback and co-design sessions. Evaluation studies often measure trust, usability, workload, perceived effectiveness, and satisfaction. Constructs such as helpfulness are typically measured with domain experts, while AI experts are generally involved in studies measuring task performance, usability, and confidence.
Table 8. Expl. Modalities
Modality Count
@@ 4.2.3. Interactivity
Table 10. Top 14 Eval. Constructs
Eval. Constructs Count
@@
Perceived Effec. 14
Task Perform. 14
@@
4.3.1. Evaluation Methods and Constructs in XUI Research Table 10
@@
categorises the various study methods used to evaluate XUIs. Controlled experiments are the most frequent ones, highlighting the importance of rigorous, systematic evaluation under controlled conditions. This method allows researchers to isolate specific factors and measure their impact on user interaction with XUIs. Additionally, interviews and usability studies are frequently utilised, highlighting the focus on user perspectives and ensuring that the interfaces remain practical and accessible.
@@
Moving onto the evaluation constructs, illustrated in Table 10, trust stands out as central, indicating that fostering user confidence in AI systems is a key priority for XUI developers. Usability and workload are close behind, emphasizing the importance of ensuring the interface is easy to use and does not overwhelm the user. Perceived effectiveness and satisfaction suggest an interest in understanding users’ perceptions of how well the system supports their needs. Table 11 shows the most popular evaluation constructs. Collectively, these constructs underscore the importance of designing XUIs that not only enhance explainability but also encourage practical adoption in real-world applications. XUIs should, in fact, be designed to inspire trust, remain user-friendly, and avoid overwhelming users with excessive complexity.
@@
It is worth noting that, despite the importance of transparency in explainability research, it has only been explicitly assessed in eight studies, suggesting that current research in XUIs should place greater emphasis on incorporating transparency constructs into evaluations.
@@
Table 11. Papers that report user studies evaluating Trust, Usability, and Workload
Evaluation Constructs N. Papers References
@@
Trust 28 (Weitz et al., 2021; Kim et al., 2023a; Wysocki et al., 2023; Guo et al., 2022; Khurana et al., 2021; Lee et al., 2020; Panigutti et al., 2023; Lu et al., 2024; Karran et al., 2022; Bhattacharya et al., 2023; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Bhattacharya et al., 2024; Kim et al., 2023d; Röhrl et al., 2023; Guesmi et al., 2022; Joshi et al., 2024; Lombardi et al., 2024; Naiseh et al., 2023; Bahel et al., 2024; Guesmi et al., 2024; Lee et al., 2023; Perlmutter et al., 2024; Bertrand et al., 2023a; Park et al., 2022; Zytek et al., 2022; Purificato et al., 2023; Heimerl et al., 2022; Nimmo et al., 2024; Lei et al., 2024; Desolda et al., 2023; Stites et al., 2021)
@@
Usability 24 (Fujiwara et al., 2020; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Wang et al., 2021; Huang et al., 2022; Cheng et al., 2021; Kim et al., 2023b; Szymanski et al., 2024; Hohman et al., 2019; Guesmi et al., 2024; Wang et al., 2022; Ming et al., 2019; Sevastjanova et al., 2021; Park et al., 2022; Zhang et al., 2024; El-Zanfaly et al., 2023; Sciascio et al., 2017; Purificato et al., 2023; Hao et al., 2023; Yuan et al., 2024b; Antar et al., 2024; Szymanski et al., 2021)
@@
Workload 15 (Fujiwara et al., 2020; Piorkowski et al., 2023; Dominguez et al., 2020; Sanneman and Shah, 2022a; Lee et al., 2020; Wang et al., 2023; Karran et al., 2022; Khanna et al., 2022; Anderson et al., 2020; El-Zanfaly et al., 2023; Kridalukmana et al., 2022; Sciascio et al., 2017; Nakao et al., 2022; Heimerl et al., 2022; Yuan et al., 2024a)
@@
Figure 5. Methods used to evaluate explanation interfaces across different user types. The heatmap shows the distribution of evaluation methods, categorised by the type of user involved (e.g., AI experts, domain experts, non-experts).
@@
4.3.2. Tailoring XUI Evaluation Methods to User Groups
Figure 5 represents the distribution of user types across various study methodologies used in the evaluation of XUIs. The figure provides some insights into how different user groups are targeted and about the methodological approaches used to assess their interaction with and understanding of the interface. The choice of study method is influenced by the target user group: controlled experiments are favoured for domain experts and lay users due to their rigorous structure. Interviews, however, are particularly prominent in studies involving domain experts, as they can usually capture qualitative insights. Domain experts are typically involved in interviews for the application of XUI technologies in the high stakes domains such as Health (Hwang et al., 2022b; Kim et al., 2023a, c; Bhattacharya et al., 2023, 2024; Naiseh et al., 2023; Okolo et al., 2024; Larasati et al., 2023; Lindvall et al., 2021; Zhang et al., 2024; Cheng et al., 2022) and Finance and economics, (Füßl et al., 2024; Purificato et al., 2023) but also in other domain such as Education (Lu et al., 2024), Natural Language Processing (Wang et al., 2022) and weather forecasting (Kim et al., 2023d). Methods such as interactive feedback sessions, co-design, usability studies, and, notably, surveys are also employed to evaluate AI experts in various domains, such as health (Hohman et al., 2019) associated with Artificial Intelligence (Baniecki et al., 2024), education (Shin et al., 2022; Yuan et al., 2024b) and finance and economics (Hohman et al., 2019; Meza Martínez et al., 2023). While these approaches are often associated with assessing non-expert users, their application to AI experts suggests an interest in gathering qualitative and usability-focused insights from technically proficient individuals.